# Scheme-independent determination of the QCD running coupling at all scales from jet observables using the Principle of Maximum Conformality/Infinity

### v2

@@ Abstract
We present a new approach to determining the strong coupling $\alpha_{s}(Q)$, over the entire range of validity of perturbative QCD, for scales above $\Lambda_{\mathrm{QCD}}$ and up to the Planck scale $\sim 1.22\cdot 10^{19}$ GeV, with the highest precision and using the data of a single experiment. In particular, we use the results obtained for the thrust ($T$) and $C$-parameter ($C$) distributions in $e^{+}e^{-}$ annihilation at a single annihilation energy $\sqrt{s}=M_{Z}$ (i.e. at the $Z^{0}$ peak). This new method is based on the intrinsic conformality (iCF) and on the Infinite-Order Scale Setting, using the Principle of Maximum Conformality (i.e. the PMC), which allows a rigorous determination of the renormalization scales for the event-shape variable distributions satisfying all of the requirements of Renormalization Group Invariance, including renormalization-scheme independence and consistency with Abelian theory in the $N_{C}\to 0$ limit. This new method is based on the scale-invariance of the iCF, which allows determination of $\alpha_{s}(\mu_{0})$ at any scale $\mu_{0}$, and on the Maximum Likelihood statistical approach. We introduce a novel approach to determining the best-fitting range based on the most-likely-lowest $\chi^{2}$ calculated considering all possible intervals among the entire range of bins available in the perturbative region. This new method improves the precision and leads to results that are statistically more reliable. In particular, we obtain the following results for the running coupling: $\alpha_{s}(M_{Z})=0.1170^{+0.0012}_{-0.0010}$ from thrust and $\alpha_{s}(M_{Z})=0.1181^{+0.0013}_{-0.0009}$ for $C$-parameter in the $\overline{\mathrm{MS}}$ scheme at NNLO. These determinations of $\alpha_{s}(M_{Z})$ are consistent with the world average and are more precise than the values obtained from the analysis of event shape observables currently used in the world average.
@@ II Event-shape variables and the renormalization scale
where $\vec{p}_{i}$ indicates the three-momentum of particle $i$ and the sums run over all particles in the hadronic final state; $\vec{n}$ is a unit vector that is varied to maximize thrust $T$, determining the so-called thrust axis, i.e. $\vec{n}_{T}$. Alternatively, the variable $(1{-}T)$ is often used, which for 3-jet production at LO is restricted to the range $0<(1{-}T)<\frac{1}{3}$. We have back-to-back and spherically symmetric events for $T=1$ and $T=\frac{2}{3}$ respectively. For the $C$-parameter, $\theta_{ij}$ is the angle between $\vec{p_{i}}$ and $\vec{p_{j}}$. For 3-jet production at LO the $C$-parameter is restricted by kinematics to the range $0\leq C\leq\frac{3}{4}$. Theoretical results for normalized infrared-safe observables, such as the thrust or $C$-parameter distribution for the annihilation process $e^{+}e^{-}\rightarrow 3\,$jets at the $Z^{0}$ peak to NNLO accuracy, are given in Refs. [50, 51, 33, 34, 35, 36, 37]. When using the conventional scale-setting method, only one value of $\alpha_{s}$, at the kinematic scale $\sqrt{s}$ of the process, can be extracted with errors given by the uncertainty in choosing the renormalization scale, while using the PMC, which is based on intrinsic conformality (iCF), it is possible to determine the strong coupling at any scale, even up to the highest scale where the theory is expected to breakdown. As shown in Eq. (I), the final renormalization scale $\mu$ is set by using the PMC method while $\mu_{0}$ can be set to any value in the perturbative range; we can thus determine $\alpha_{s}(\mu_{0})$ at any scale. Results for the PMC for both thrust and $C$-parameter have been discussed in Refs. [2, 52, 3, 53]. We remark that the PMC solves the renormalization scale ambiguity and the scales, related to each order of calculation, are fixed perturbatively by determining the $\beta_{0}$ term of each scale at each order of accuracy. These scales take into account the dynamics of the underlying quark and gluon subprocess and thus fix the magnitude of the strong coupling at each order and for the corresponding value of the selected variable $(C,1{-}T))$. In fact, these scales are not single valued but are themselves functions of the kinematic scale, $\sqrt{s}$, and of the event-shape variable via the coefficients of the number of flavors, $N_{f}$, related to the loop integrals and thus to their UV-divergent part. The PMC scales are fixed order by order and the last scale is set to the kinematic scale of the process $\sqrt{s}\equiv M_{Z}$, given the missing knowledge of the higher-order $\beta_{0}$-term. We have defined this type of uncertainty as the second kind of residual scale dependence [54]. We remark that both the first and second kind of residual scale uncertainties are of the same order as the uncertainties in the first missing higher-order contribution and thus both have a mere perturbative nature linked to the perturbative series and its accuracy.
@@ III Fitting Method
The $\chi^{2}$ analysis has been performed following the approach of Ref. [55]; thus, in the first approximation, considering the bins as being independent of each other and the experimental errors $\sigma_{i}$ determined by summing the statistic and systematic errors (due to detector effects and to hadronization) in quadrature. However, bin-to-bin correlation effects are introduced when evaluating systematic uncertainties and thus the errors on the fitted parameter and the same $\chi_{\mathrm{min}}^{2}$ values are considered only for sake of comparison and not as absolute results.
@@ IV Results
with the same statistics $\chi^{2}/n_{\mathrm{dof}}$ and $P_{\chi^{2}}(\chi^{2}_{\mathrm{min}};n_{\mathrm{dof}})$, i.e. $0.8/11$ and $1.6\cdot 10^{-5}$ respectively. These results lead to the same predictions for $\alpha_{s}(M_{Z})$ shown in Table 2. Analogous results are obtained in the other bin ranges. We note that the results shown in Table 2, for thrust using the PMC and the 5-loop solution for the running coupling have very small differences with respect to the particular range and that the same prediction is obtained for the coupling using the 5-loop perturbative solution of Eq. (I), as shown in Table 3 in the BFR range. It is to be noted that in the case of thrust results for extracted values of the coupling using the PMC in the BFR and in the total range of bins (TFR) are consistent within the errors. While for the $C$-parameter we find a significant reduction of the errors, of one order of magnitude in the range TFR1 using the PMC, independently from the particular solution used for the coupling as shown in Table 5 and Table 3. This effect is possibly due to an effect of constraint of the particular bin at $C\sim 0.05$. In fact, in the range of bins TFR2 that excludes this particular bin, the effect disappear. Given the poor statistics, we would tend to exclude this prediction, but in a global fit that includes also nonperturbative effects over the entire range of values (TFR1) of the $C$-parameter and that would improve the goodness of the fit in this range, we would expect a higher precision in the determination of the strong coupling. We show results using the conventional scale setting for thrust and the $C$-parameter in Table 4 and Table 7 respectively. Comparing the results to those shown in Ref. [4], we note that, by using the conventional scale setting, we obtain the same errors for the strong coupling extracted from the NNLO theoretical predictions of thrust and the $C$-parameter but with improved statistics. The extracted values of the strong coupling are analogous to those of Ref. [4]; although in the case of the $C$-parameter we find a small deviation. This is possibly due to the small differences in the NNLO coefficient that have also been shown in Ref. [37]. Comparison between conventional scale setting and PMC results clearly shows that PMC leads to a significantly higher precision.
@@
$\alpha_{s}(M_{Z})=0.1181^{+0.0005}_{-0.0005},$ (10)
which is particularly consistent with the world average; i.e. $\alpha_{s}(M_{Z})=0.1180\pm 0.0009$ [1].
@@
In Fig. 3 we show the strong coupling $\alpha_{s}(\mu)$ behavior determined by the weighted average and also the results for $\alpha_{s}(M_{Z})$ from thrust and the $C$-parameter. Results for both thrust and the $C$-parameter are consistent with the world average of $\alpha_{s}(M_{Z})=0.1180\pm 0.0009$ [1] and also with recent PMC applications presented in Refs. [57, 58, 59, 60].

### v3

# Scheme-independent determination of the QCD running coupling at all scales from jet observables using the principle of maximum conformality and infinite-order scale setting

@@ Abstract
We present a new approach to determining the strong coupling $\alpha_{s}(Q)$, over the entire range of validity of perturbative QCD, for scales above $\Lambda_{\mathrm{QCD}}$ and up to the Planck scale $\sim 1.22\cdot 10^{19}$ GeV, with the highest precision and using the data of a single experiment. In particular, we use the results obtained for the thrust ($T$) and $C$-parameter ($C$) distributions in $e^{+}e^{-}$ annihilation at a single annihilation energy $\sqrt{s}=M_{Z}$ (i.e. at the $Z^{0}$ peak). This new method is based on the intrinsic conformality (iCF) and on the Infinite-Order Scale Setting, using the Principle of Maximum Conformality (i.e. the PMC), which allows a rigorous determination of the renormalization scales for the event-shape variable distributions satisfying all of the requirements of Renormalization Group Invariance, including renormalization-scheme independence and consistency with Abelian theory in the $N_{C}\to 0$ limit. This new method is based on the scale-invariance of the iCF, which allows determination of $\alpha_{s}(\mu_{0})$ at any scale $\mu_{0}$, and on the Maximum Likelihood statistical approach. We introduce a novel approach to determining the best-fitting range based on the most-likely-lowest $\chi^{2}$ calculated considering all possible intervals among the entire range of bins available in the perturbative region. This new method improves the precision and leads to results that are statistically more reliable. In particular, using data for thrust and $C$-parameter at the $Z^{0}$ peak from ALEPH, OPAL, DELPHI and L3 experiments, we obtain the average value: $\alpha_{s}(M_{Z})=0.1181^{+0.0005}_{-0.0005}$, for the strong coupling. This determination of $\alpha_{s}(M_{Z})$ is consistent with the world average and has an improved precision with respect to the values obtained from the analysis of event shape observables currently used in the world average.
@@ II Event-shape variables and the renormalization scale
where $\vec{p}_{i}$ indicates the three-momentum of particle $i$ and the sums run over all particles in the hadronic final state; $\vec{n}$ is a unit vector that is varied to maximize thrust $T$, determining the so-called thrust axis, i.e. $\vec{n}_{T}$. Alternatively, the variable $(1{-}T)$ is often used, which for 3-jet production at LO is restricted to the range $0<(1{-}T)<\frac{1}{3}$. We have back-to-back and spherically symmetric events for $T=1$ and $T=\frac{2}{3}$ respectively. For the $C$-parameter, $\theta_{ij}$ is the angle between $\vec{p_{i}}$ and $\vec{p_{j}}$. For 3-jet production at LO the $C$-parameter is restricted by kinematics to the range $0\leq C\leq\frac{3}{4}$.
Theoretical results for normalized infrared-safe observables, such as the thrust or $C$-parameter distribution for the annihilation process $e^{+}e^{-}\rightarrow 3\,$jets at the $Z^{0}$ peak to NNLO accuracy, are given in Refs. [50, 51, 33, 34, 35, 36, 37]. When using the conventional scale-setting method, only one value of $\alpha_{s}$, at the kinematic scale $\sqrt{s}$ of the process, can be extracted with errors given by the uncertainty in choosing the renormalization scale, while using the PMC, which is based on intrinsic conformality (iCF), it is possible to determine the strong coupling at any scale, even up to the highest scale where the theory is expected to breakdown. As shown in Eq. (I), the final renormalization scale $\mu$ is set by using the PMC method while $\mu_{0}$ can be set to any value in the perturbative range; we can thus determine $\alpha_{s}(\mu_{0})$ at any scale. Results for the PMC for both thrust and $C$-parameter have been discussed in Refs. [2, 52, 3, 53]. We remark that the PMC solves the renormalization scale ambiguity and the scales, related to each order of calculation, are fixed perturbatively by determining the $\beta_{0}$ term of each scale at each order of accuracy. These scales take into account the dynamics of the underlying quark and gluon subprocess and thus fix the magnitude of the strong coupling at each order and for the corresponding value of the selected variable $(C,1{-}T))$. In fact, these scales are not single valued but are themselves functions of the kinematic scale, $\sqrt{s}$, and of the event-shape variable via the coefficients of the number of flavors, $N_{f}$, related to the loop integrals and thus to their UV-divergent part. The PMC scales are fixed order by order and the last scale is set to the kinematic scale of the process $\sqrt{s}\equiv M_{Z}$, given the missing knowledge of the higher-order $\beta_{0}$-term. We have defined this type of uncertainty as the second kind of residual scale dependence [54]. We remark that both the first and second kind of residual scale uncertainties are of the same order as the uncertainties in the first missing higher-order contribution and thus both have a mere perturbative nature linked to the perturbative series and its accuracy.
@@ III Fitting Method
The $\chi^{2}$ analysis has been performed following the approach of Ref. [55]; thus, in the first approximation, considering the bins as being independent of each other and the experimental errors $\sigma_{i}$ determined by summing the statistic and systematic errors (due to detector effects and to hadronization) in quadrature. Thus, we have not considered correlation effects and the errors calculated on the fitted parameter and the same $\chi_{\mathrm{min}}^{2}$-values are considered mainly for sake of comparison among different fitting-regions, different models for the coupling and different methods for the scale-setting; only secondarily these values have been used to determine the goodness-of-fit. In fact, as reported in Ref. [55], without the inclusion of the correlation effects these values can only give a rough measure for the goodness-of-fit.
@@ IV Results
with the same statistics $\chi^{2}/n_{\mathrm{dof}}$ and $P_{\chi^{2}}(\chi^{2}_{\mathrm{min}};n_{\mathrm{dof}})$, i.e. $0.8/11$ and $1.6\cdot 10^{-5}$ respectively. These results lead to the same predictions for $\alpha_{s}(M_{Z})$ shown in Table 2. We remark that in this approach, the physical scale is given by the final renormalization scale $\mu_{\rm PMC}$ set by using the PMC method, while the initial scale, $\mu_{0}$, is purely conventional and is related to the final renormalization scale by renormalization group transformations. Thus, the determination of the coupling at the Planck scale $\alpha_{s}(M_{\rm Pl})$, should not be considered as a direct measure of the coupling at the physical scale $M_{\rm Pl}$, but only as a test of the PMC method.
Analogous results are obtained in the other bin ranges. We note that the results shown in Table 2, for thrust using the PMC and the 5-loop solution for the running coupling have very small differences with respect to the particular range and that the same prediction is obtained for the coupling using the 5-loop perturbative solution of Eq. (I), as shown in Table 3 in the BFR range. It is to be noted that in the case of thrust results for extracted values of the coupling using the PMC in the BFR and in the total range of bins (TFR) are consistent within the errors. While for the $C$-parameter we find a significant reduction of the errors, of one order of magnitude in the range TFR1 using the PMC, independently from the particular solution used for the coupling as shown in Table 5 and Table 3. This effect is possibly due to an effect of constraint of the particular bin at $C\sim 0.05$. In fact, in the range of bins TFR2 that excludes this particular bin, the effect disappear. Given the poor statistics, we would tend to exclude this prediction, but in a global fit that includes also nonperturbative effects over the entire range of values (TFR1) of the $C$-parameter and that would improve the goodness of the fit in this range, we would expect a higher precision in the determination of the strong coupling. We show results using the conventional scale setting for thrust and the $C$-parameter in Table 4 and Table 7 respectively. Comparing the results to those shown in Ref. [4], we note that, by using the conventional scale setting, we obtain the same errors for the strong coupling extracted from the NNLO theoretical predictions of thrust and the $C$-parameter but with improved statistics. The extracted values of the strong coupling are analogous to those of Ref. [4]; although in the case of the $C$-parameter we find a small deviation. This is possibly due to the small differences in the NNLO coefficient that have also been shown in Ref. [37]. Comparison between conventional scale setting and PMC results clearly shows that PMC leads to a significantly higher precision.
@@
$\alpha_{s}(M_{Z})=0.1181^{+0.0005}_{-0.0005}.$ (10)
@@
In Fig. 3 we show the strong coupling $\alpha_{s}(M_{Z})$ result determined by the weighted average and also the results for $\alpha_{s}(M_{Z})$ from thrust and the $C$-parameter. Results for both thrust and the $C$-parameter are consistent with the world average [1] and also with recent PMC applications presented in Refs. [57, 58, 59, 60]. We point out that in this analysis we have not included part of the systematic errors, in particular those related to the hadronization effects. Thus, a direct comparison of the total error in Eq.10, with the total error obtained by summing in quadrature the statistical and theoretical errors (not the systematic error) obtained in Ref. [4], (i.e. $\Delta\alpha_{s}(M_{Z})\simeq\pm 0.0036$), shows that the PMC method can lead to a clear improvement of the precision with respect to the conventional scale setting adopted in Ref. [4].